OpenStudy (anonymous):

om's and Tim's ages add up to 21 years. In 3 years, Tom will be twice Tim’s age. How old are they now?

OpenStudy (uri):

2x+x=21 3x=21 x=21/3 x=7

OpenStudy (uri):

Tim =7 If you want to know how old is Tom 2(7) =14

OpenStudy (uri):

@zxcvbnm08765

OpenStudy (anonymous):

but @uri they said in 3 years,tom will be twice tim's age

OpenStudy (uri):

Oh.

OpenStudy (anonymous):

suppose we add 21+3=24 right?

OpenStudy (uri):

You're right.

OpenStudy (anonymous):



OpenStudy (anonymous):